OpenStudy (anonymous):

Your Math class has 32 students. You would like to treat the class to cookies. You would like each person to have 5 cookies. Which expression would be a quick way to figure out how many cookies you would need to bake?

OpenStudy (anonymous):

if there are 32 students eating five cookies, then you would have 32 * 5 cookies total cookies = 32 students * 5 cookies /student

OpenStudy (anonymous):

heres the possible choices 5(32 + 2) 5(30 + 2) 5(32 − 2) 5(30 − 2)

OpenStudy (anonymous):

32 is the same as 30 + 2 so 5 * 32 = 5(30 + 2)

OpenStudy (anonymous):

this has to be the strangest question i have seen today i would multiply 32 by 5

OpenStudy (anonymous):

it is not an equation, it is a multiplication

OpenStudy (anonymous):

how on earth did you find that????

OpenStudy (saifoo.khan):

One more: http://openstudy.com/study#/updates/4f9eee68e4b000ae9ed329c0

OpenStudy (anonymous):

is that some idiot math teachers idea of finding a quick way to multiply 32 by 5? how about half of 32 is 16, add a zero get 160

OpenStudy (anonymous):

all right show off cut it out and tell me your secret

OpenStudy (saifoo.khan):

LOL^

OpenStudy (anonymous):

u just have to look on google

OpenStudy (saifoo.khan):

Don't just look. Type it too. xD

OpenStudy (anonymous):

scary